OpenStudy (precal):

A differentiable function f on the interval -2

OpenStudy (precal):

A. f(7) is the absolute maximum value of f on the interval [-2,7] B. f(c) is the absolute minimum value of f on the interval [-2,7] C. Either f(-2) or f(2) is the absolute minimum on the interval [-2,7] D. Either f(c) or f(7) is the absolute minimum on the interval [-2,7] E. The value of f ' (x)<0 on the interval [c,7]

OpenStudy (precal):

I know the solution is C but I want to know why

OpenStudy (anonymous):

So \(f\) is differentiable on \([-2,7]\) (this should be an open interval; a function can't be differentiable at an endpoint - in any case, we know the function is continuous) and \(f'(c)=0\) for some \(-2<c<7\). Do you recall the mean value theorem?

OpenStudy (precal):

yes where f ' (c) = f(7)-f(-2) over 9

OpenStudy (precal):

in this case

OpenStudy (anonymous):

Ooh, actually Rolle's theorem is the one I meant to say

OpenStudy (anonymous):

Rolle's is just a specific case of the MVT, at any rate. We know that for this \(c\), we have \[\frac{f(7)-f(-2)}{9}=0~~\iff~~f(7)=f(-2)\]

OpenStudy (precal):

thanks I can't seem to get those thm straight

OpenStudy (anonymous):

No problem. I don't see how C could be right though. How could we say anything definitive about \(f(2)\)?

OpenStudy (precal):

who knows? Could be a typo on the part of the author

OpenStudy (anonymous):

Hmm, I would have gone with B.